OpenStudy (anonymous):

Look at the table of values below. a b 2 8 4 14 6 20 Which equation is represented by the table?

OpenStudy (anonymous):

b = a+14/2 b = 3a + 2 b = 4a - 4 b = 4a - 2

OpenStudy (anonymous):

@Data_LG2

OpenStudy (anonymous):

here's the tip, try all the values into the choices to see if they satisfy one of those.\ for example you have (a,b)=(2,8) substitute it ont he first equation which is b = a+14/2 so 8=2 + 14/2 8=2+7 8=9 -> not equal it means this is not the right one :) try it on the other choices.

OpenStudy (anonymous):

Do you know this @Klip

OpenStudy (anonymous):

Help plz

OpenStudy (anonymous):

Well, now you know it's not A. Try B, b=3a+2 (a,b)=(2,8) 8=3(2)+2 8=6+2 8=8

OpenStudy (anonymous):

There you go

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

You can give medals by clicking the Best Response button.